The sum of a number and 15 is 63. Find the number ?



Answer by lenny460(1073)

The sum of a number and 15 is 63. Find the number ?
Let the number be x.
x + 15 = 63
Subtract 15 from each side of the equation.
x + 15 - 15 = 63 - 15
Combine like terms:
15 - 15 = 0
63 - 15 = 48
x = 48


Answer:
The number is 48







lennox obuong
algebra tutor